1.     What happens to the speed, does it slow down as different objects are added and the applied force is different?  Why do you think this happens?

 

2.     If forces are balanced, then the net force is zero. Is there a sum of forces, or net force here?


1
Expert's answer
2021-12-27T08:00:59-0500

Explanation & Answer


  • The kinetic energy gain as a result of work done by a force moved some distance is given by,

"\qquad\qquad
\begin{aligned}
\small Fs&=\small \frac{1}{2}mv^2
\end{aligned}"

  • Then the speed change is due to the change of force and mass for a same distance moved every time is obtained as follows.

"\qquad\qquad
\begin{aligned}
\small v&=\small \sqrt{\frac{2Fs}{m}}\\
&\propto \sqrt{\frac{F}{m}}
\end{aligned}"

  • The speed will decrease if the mass was raised while keeping the force constant and vise versa.



  • When the net force is zero, there are forces acting on the body and as a result of that, the net force or the resulting force turns out to be zero.
  • In other words, there is a sum of forces and it is zero when the forces are balanced.